courtneyd11:

on a cloudy day, young son needed to know the height of a window in a building, young son, position a mirror on the ground between herself and the building so that when he looked in the mirror, he saw the window if the mirror was 22.88 cm from her feet and 34.15 m from the base of the building and young sons I was 1.89 m above the ground how high up on the building was the window located

Alexis1415:

@ihatelifefs420 wrote:
22.88cm/1.89m=34.15m/x solving it for x , we get x=(34.14*1.89)/0.2288 (we have converted 22.88cm into meters) =282m so 282m is the answer
Did you get this from a site.......like.....please do not plagiarize fr. Next time do your own work please and thank you.

ihatelifeFS420:

@alexis1415 wrote:
@ihatelifefs420 wrote:
22.88cm/1.89m=34.15m/x solving it for x , we get x=(34.14*1.89)/0.2288 (we have converted 22.88cm into meters) =282m so 282m is the answer
Did you get this from a site.......like.....please do not plagiarize fr. Next time do your own work please and thank you.
I got it from this site somebody already answered this question on here before